OpenStudy (idku):

One end of a string 7 m long is moved up and down with simple harmonic motion at a frequency of 70 Hz. The waves reach the other end of the string in 0.4 s. Find the wavelength of the waves on the string.

OpenStudy (perl):

Hint: frequency*wavelength=speed

OpenStudy (idku):

\(\displaystyle v=\frac{L}{t}=\frac{7}{0.4}=17.5\) \(\displaystyle f=70\) \(\displaystyle f\times \lambda=v\) \(\displaystyle 70\times \lambda=17.5\) \(\displaystyle 70\times \lambda=0.25~{\rm meters}\)

OpenStudy (idku):

yeah, I saw "frequency*wavelength=speed" the first time I looked online (our teacher didn't cover this yet assigned for h/w :( ), so I was thinking about some way of putting it.

OpenStudy (idku):

Is this correct or totally ridiculous ?

OpenStudy (idku):

Oh, yes, it's correct;) Thanks!